# Federated Learning with Reduced Information Leakage and Computation

### v1

@@ 4 Convergence Analysis
In this section, we analyze the convergence property of Upcycled-FL. Note that we do not require local functions $F_{i}(\cdot)$ to be convex. Moreover, we consider practical settings where data are non-i.i.d across different clients. Similar to [2], we introduce a measure below to quantify the dissimilarity between clients in the federated network.
Definition 4.1 ($B$-Dissimilarity [2]).
@@
where $\mathbb{E}[\cdot]$ denotes the expectation over clients (Eqn. (1)). Parameter $B\geq 1$ captures the statistical heterogeneity across different clients: when all clients are homogeneous with i.i.d. data, we have $B=1$ for all local functions; the larger value of $B$, the more dissimilarity among clients.
@@
The above assumptions are fairly standard. They first appeared in [2] and are adopted in subsequent works such as [4, 32, 33].
Note that strongly convex assumption is not on local objective $F_{i}(\omega;\mathcal{D}_{i})$, but the regularized function $F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$, i.e., the assumption can be satisfied by selecting a sufficiently large $\mu>0$. Indeed, as shown in Section 6, our algorithm still converges even when Assumption 4.2 and 4.4 do not hold (e.g., DNN). Next, we conduct the theoretical analysis on the convergence (rate) of Upcycled-FL.
Algorithm 2 Proposed framework: Upcycled-FL
2:  for $m=1$ to $M$ do
3:     The central server sends the current global model parameter $\overline{\omega}^{2m-2}$ to all the clients.
4:     A subset of clients are selected to be active and each active client updates its local model by finding (approximate) minimizer of local loss function:
$\omega_{i}^{2m-1}\leftarrow\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}.$
5:     Clients send local models to the central server.
$\overline{\omega}^{2m}$ $=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
7:  end for
@@
Let $\mathcal{S}_{m}$ be a set of $K$ randomly selected clients got updated (i.e., active clients) at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of clients. Then under Assumption 4.2 and 4.4, we have
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\overline{\omega}^{2m-1})-\boldsymbol{C_{1}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
$+\boldsymbol{C_{2}}\cdot h^{1}_{m}+\boldsymbol{C_{3}}\cdot h^{2}_{m},$
@@
The details of the term $\boldsymbol{C_{1}}$, $\boldsymbol{C_{2}}$, $\boldsymbol{C_{3}}$ (expressed as functions of $L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\lambda_{m}}$) are in Appendix E, Eqn. (10)-(12).
Lemma 4.5 characterizes the relation of the values of global objective function over two consecutive odd iterations. It is easy to verify $\boldsymbol{C_{2}},\boldsymbol{C_{3}}\geq 0$. By rearranging and telescoping, we get the following convergence rate of Upcycled-FL.
Theorem 4.6 (Convergence rate of Upcycled-FL).
Under Assumption 4.2 and 4.4, if $\boldsymbol{C_{1}}>0$, we have
$\min_{m\in[M]}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
$\leq$ $\frac{1}{M}\sum_{m=0}^{M}\mathbb{E}[||\nabla f(\overline{\omega}^{2m-1})||^{2}]$
@@
where $\overline{\omega}^{0}$ and $\overline{\omega}^{*}$ denote the initial and the optimal global model parameters, and $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ are decreasing in $\frac{\lambda_{m}}{\mu}$.
Theorem 4.6 implies that tunable $\mu$ and $\lambda_{m}$ are key hyper-parameters that control the convergence (rate) and robustness of Upcycled-FL. Recall that $\mu$ penalizes the deviation of local model $\omega_{i}^{2m}$ from global aggregated model $\overline{\omega}^{2m-1}$, while $\lambda_{m}$ penalizes the update of local model $\omega_{i}^{2m}$ from its previous update $\omega_{i}^{2m-1}$.
@@
It is worth noting that the convergence rate also depends on data heterogeneity, captured by dissimilarity $B$. According to Eqn. (10), $\boldsymbol{C_{1}}>0$ must hold when $B=0$ (i.i.d. clients). Although $\boldsymbol{C_{1}}$ may become negative as $B$ increases, the experiments in Section 6 show that Upcycled-FL can still converge when data is highly heterogeneous.
@@
Assumption 4.7 is standard and has been used when proving the convergence of FL algorithms [34, 35]; it requires that the difference of aggregated weights between two consecutive iterations and the gradient $||\nabla f(\omega)||$ are bounded. Under this assumption, we have the following corollary.
@@
Under Assumption 4.2, 4.4, and 4.7, for fixed $\mu,K$, if $\lambda_{m}$ is taken such that $\frac{\mu}{\mu+\lambda_{m}}=\mathcal{O}\Big{(}\frac{1}{\sqrt{M}}\Big{)}$, then the convergence rate of Upcycled-FL reduces to $\mathcal{O}(\frac{1}{\sqrt{M}})$.
Corollary 4.8 provides guidance on how to choose the value of $\lambda_{m}$ to guarantee the convergence of Upcycled-FL, i.e., by taking an increasing sequence of $\{\lambda_{m}\}_{m=1}^{M}$. Intuitively, increasing $\lambda_{m}$ during the training helps stabilize the algorithm, because the deviation of local models from the previous update gets penalized more under a larger $\lambda_{m}$.
@@ 5 Private Upcycled-FL
In this section, we present a privacy-preserving version of Upcycled-FL. Many perturbation mechanisms can be adopted to achieve differential privacy such as objective perturbation [36, 37], output perturbation [36, 38], gradient perturbation [39, 40], etc. Next, we will use output and objective perturbation as examples to illustrate that Upcycled-FL, by reducing the total information leakage, has a better privacy-accuracy trade-off than other methods such as FedProx, FedAvg, etc. Note that both output and objective perturbation methods are used for generating private updates at odd iterations, which can be used directly for even updates.
@@
Objective perturbation: a random linear term $\langle n^{m}_{i},\omega\rangle$ is added to the objective function in odd $(2m+1)$-th iteration, and the private local model $\widehat{\omega}_{i}^{2m+1}$ is found by solving a perturbed optimization:
@@
Lemma 5.1 is straightforward; it can be derived directly by leveraging a property of differential privacy called immunity to post-processing [41], i.e., a differentially private output followed by any data-independent computation remains satisfying differential privacy.
Based on Lemma 5.1, we can quantify the total privacy loss of private Upcycled-FL. We shall adopt moments accountant method [27] for output perturbation, and the analysis method in [36, 42, 43] for objective perturbation. These theoretical results are presented in Appendix C, Theorem C.1-C.2.
@@ 6 Experiments
In this section, we empirically evaluate Upcycled-FL and compare it with two algorithms: FedAvg [44] and FedProx [2]. We first consider non-private algorithms to examine the convergence (rate) and robustness of Upcycled-FL against statistical/device heterogeneity. Then, we adopt both output and objective perturbation to evaluate the private Upcycled-FL.
Experimental setup. We run experiments on both synthetic and real data.
@@
  • Synthetic data: using the method in [2], we generate Synthetic(iid), Synthetic(0,0), Synthetic(0.5,0.5), Synthetic(1,1), four synthetic datasets with increasing statistical heterogeneity.
  • Real data: we adopt two real datasets: 1) Femnist, a federated version of Emnist [45]; 2) Sentiment140(Sent140), a text sentiment analysis task on tweets [46]. Both of the datasets can be obtained from LEAF [47].
@@
To simulate device heterogeneity, we randomly select a fraction of devices to train at each round, and assume there are stragglers that cannot train for full rounds; both devices and stragglers are selected by random seed to ensure they are the same for all algorithms. We also consider several learning algorithms: we learn logistic regression on synthetic data; multilayer perceptron (MLP) and deep CNN on Femnist; Stacked LSTM on Sent140. For Femnist, we present the results on MLP here and other results in Appendix H.
Because even iterations of Upcycled-FL only involve addition/subtraction operations with almost no computational cost, we train Upcycled-FL with double iterations compared to FedAvg and FedProx in approximately same training time. Unless explicitly stated, the results we report are averaged outcomes over all devices. More details of experimental setup are in Appendix H.1.
(a) Synthetic(0.5,0.5)
(b) Femnist
Figure 2: Convergence of Upcycled-FL, FedProx, FedAvg under the approximate same training time (left) and impact of $\lambda_{m}$ on convergence rate of Upcycled-FL (right).
Table 1: Training(testing) accuracy of FedAvg, FedProx, Upcycled-FL with 90% straggler: FedAvg/FedProx/Upcycled-FL are trained over synthetic data for 80/80/160 iterations, and are trained over Femnist for 150/150/300 iterations. For Upcycled-FL, $\lambda=0.04,0.12,0.21,0.43$ is set based on the value of $\mu=0.1,0.3,0.5,1$, respectively. Upcycled-FL significantly outperforms FedAvg and FedProx.
Dataset Proximal term Accuracy(%)
FedAvg FedProx Upcycled-FL
Synthetic(iid) $\mu=0.1$ 97.80%(97.51%) 96.86%(96.34%) 97.88%(97.51%)
$\mu=0.3$ 95.18%(94.88%) 96.31%(96.05%)
$\mu=0.5$ 94.03%(93.70%) 95.40%(95.02%)
$\mu=1$ 92.14%(91.80%) 93.76%(93.70%)
Synthetic(0,0) $\mu=0.1$ 78.41%(76.27%) 81.66%(79.50%) 84.41%(82.44%)
$\mu=0.3$ 81.73%(79.35%) 83.83%(81.58%)
$\mu=0.5$ 81.55%(79.35%) 82.84%(81.29%)
$\mu=1$ 81.22%(78.85%) 82.41%(80.57%)
Synthetic(0.5,0.5) $\mu=0.1$ 79.53%(78.97%) 79.63%(79.58%) 81.19%(81.54%)
$\mu=0.3$ 80.31%(80.56%) 82.14%(83.74%)
$\mu=0.5$ 79.55%(79.46%) 82.61%(82.64%)
$\mu=1$ 80.11%(80.20%) 82.74%(82.64%)
Synthetic(1,1) $\mu=0.1$ 69.70%(71.12%) 79.20%(78.83%) 77.38%(78.64%)
$\mu=0.3$ 74.77%(75.91%) 80.43%(81.47%)
$\mu=0.5$ 75.36%(76.67%) 82.78%(83.07%)
$\mu=1$ 75.69%(76.58%) 82.88%(82.69%)
Femnist $\mu=0.1$ 20.43%(20.01%) 82.86%(83.73%) 86.37%(85.91%)
$\mu=0.3$ 76.10%(76.66%) 81.94%(82.12%)
$\mu=0.5$ 68.28%(68.87%) 77.77%(77.60%)
$\mu=1$ 28.82%(29.05%) 54.81%(54.00%)
Sent140 $\mu=0.1$ 73.41%(71.13%) 75.29%(72.57%) 78.12%(72.38%)
$\mu=0.3$ 72.86%(71.38%) 74.27%(73.02%)
$\mu=0.5$ 70.18%(69.07%) 72.92%(71.44%)
$\mu=1$ 62.10%(62.13%) 63.91%(64.11%)

### v2

@@ 5 Convergence Analysis
In this section, we analyze the convergence of Upcycled-FL. For illustrating purposes, we focus on analyzing the convergence of Upcycled-FedProx. Note that we do not require local functions $F_{i}(\cdot)$ to be convex. Moreover, we consider practical settings where data are non-i.i.d across different clients. Similar to Li et al. (2020), we introduce a measure below to quantify the dissimilarity between clients in the federated network.
Definition 5.1 ($B$-Dissimilarity (Li et al., 2020)).
@@
where $\mathbb{E}[\cdot]$ denotes the expectation over clients (see Eqn. (1)). Parameter $B\geq 1$ captures the statistical heterogeneity across different clients: when all clients are homogeneous with i.i.d. data, we have $B=1$ for all local functions; the larger value of $B$, the more dissimilarity among clients.
@@
The above assumptions are fairly standard. They first appeared in Li et al. (2020) and are adopted in subsequent works such as T Dinh et al. (2020); Khaled et al. (2020); Pathak and Wainwright (2020). Note that strongly convex assumption is not on local objective $F_{i}(\omega;\mathcal{D}_{i})$, but the regularized function $F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$, i.e., the assumption can be satisfied by selecting a sufficiently large $\mu>0$. Indeed, as shown in Section 7, our algorithm still converges even when Assumption 5.2 and 5.4 do not hold (e.g., DNN). Next, we conduct the theoretical analysis on the convergence rate.
@@
Let $\mathcal{S}_{m}$ be a set of $K$ randomly selected clients which got updated (i.e., active clients) at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of clients. Then under Assumption 5.2 and 5.4, we have
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$$\leq$ $f(\overline{\omega}^{2m-1})-\boldsymbol{C_{1}}||\nabla f(\overline{\omega}^{2m-1})||^{2}+\boldsymbol{C_{2}}\cdot h^{1}_{m}+\boldsymbol{C_{3}}\cdot h^{2}_{m},$
@@
The details of term $\boldsymbol{C_{1}}$, $\boldsymbol{C_{2}}$, $\boldsymbol{C_{3}}$ (expressed as functions of $L,B,\frac{1}{\mu},\frac{1}{\rho},\frac{1}{K},\frac{\mu}{\lambda_{m}}$) are in Appendix C, Eqn. (11)-(13).
Lemma 5.5 characterizes the relation of the values of global objective function over two consecutive odd iterations. It is easy to verify $\boldsymbol{C_{2}},\boldsymbol{C_{3}}\geq 0$. By rearranging and telescoping, we get the following convergence rate of Upcycled-FedProx.
Theorem 5.6 (Convergence rate of Upcycled-FedProx).
Under Assumption 5.2 and 5.4, if $\boldsymbol{C_{1}}>0$, we have
$\min_{m\in[M]}\mathbb{E}\left[||\nabla f(\overline{\omega}^{2m-1})||^{2}\right]$$\leq$ $\frac{1}{M}\sum_{m=0}^{M}\mathbb{E}\left[||\nabla f(\overline{\omega}^{2m-1})||^{2}\right]$
@@
where $\overline{\omega}^{0}$ and $\overline{\omega}^{*}$ denote the initial and the optimal global model parameters, respectively. Both terms $\boldsymbol{C_{2}}$ and $\boldsymbol{C_{3}}$ are decreasing in $\frac{\lambda_{m}}{\mu}$.
Theorem 5.6 implies that tunable $\mu$ and $\lambda_{m}$ are key hyper-parameters that control the convergence (rate) and robustness of Upcycled-FedProx. Recall that $\mu$ penalizes the deviation of local model $\omega_{i}^{2m}$ from global aggregated model $\overline{\omega}^{2m-1}$, while $\lambda_{m}$ penalizes the update of local model $\omega_{i}^{2m}$ from its previous update $\omega_{i}^{2m-1}$.
@@
It is worth noting that the convergence rate also depends on data heterogeneity, captured by dissimilarity $B$. According to Eqn. (11), $\boldsymbol{C_{1}}>0$ must hold when $B=0$ (i.i.d. clients). Although $\boldsymbol{C_{1}}$ may become negative as $B$ increases, the experiments in Section 7 show that Upcycled-FedProx can still converge when data is highly heterogeneous.
@@
Assumption 5.7 is standard and has been used when proving the convergence of FL algorithms (Li et al., 2019; Yang et al., 2022); it requires that the difference of aggregated weights between two consecutive iterations and the gradient $||\nabla f(\omega)||$ are bounded. Under this assumption, we have the following corollary.
@@
Under Assumption 5.2, 5.4, and 5.7, for fixed $\mu,K$, if $\lambda_{m}$ is taken such that $\frac{\mu}{\mu+\lambda_{m}}=\mathcal{O}\Big{(}\frac{1}{\sqrt{M}}\Big{)}$, then the convergence rate of Upcycled-FedProx reduces to $\mathcal{O}(\frac{1}{\sqrt{M}})$.
Corollary 5.8 provides guidance on selecting the value of $\lambda_{m}$ properly to guarantee the convergence of Upcycled-FedProx, i.e., by taking an increasing sequence of $\{\lambda_{m}\}_{m=1}^{M}$. Intuitively, increasing $\lambda_{m}$ during the training helps stabilize the algorithm, because the deviation of local models from the previous update gets penalized more under a larger $\lambda_{m}$.
@@ 6 Private Upcycled-FL
In this section, we present a privacy-preserving version of Upcycled-FL. Many perturbation mechanisms can be adopted to achieve differential privacy such as objective perturbation (Chaudhuri et al., 2011; Kifer et al., 2012), output perturbation (Chaudhuri et al., 2011; Zhang et al., 2017), gradient perturbation (Bassily et al., 2014; Wang et al., 2017), etc. In this section, we use output and objective perturbation as examples to illustrate that FL algorithms combined with Upcycled-FL strategy, by reducing the total information leakage, can attain a better privacy-accuracy trade-off. Note that both output and objective perturbation methods are used to generate private updates at odd iterations, which can be used directly for even updates.
@@
Objective perturbation: a random linear term $\langle n^{m}_{i},\omega\rangle$ is added to the objective function in odd $(2m+1)$-th iteration, and the private local model $\widehat{\omega}_{i}^{2m+1}$ is found by solving a perturbed optimization.
Taking Upcycled-FedProx as the example, we have:
@@
Lemma 6.1 is straightforward; it can be derived directly by leveraging a property of differential privacy called immunity to post-processing (Dwork et al., 2014), i.e., a differentially private output followed by any data-independent computation remains satisfying differential privacy.
Based on Lemma 6.1, we can quantify the total privacy loss of private Upcycled-FL. We shall adopt moments accountant method (Abadi et al., 2016) for output perturbation, and the analysis method in (Chaudhuri et al., 2011; Zhang and Zhu, 2016; Zhang et al., 2018b) for objective perturbation.
Here, we focus on settings where local loss function $F_{i}(\omega_{i},\mathcal{D}_{i}):=\frac{1}{|\mathcal{D}_{i}|}\sum_{d\in\mathcal{D}_{i}}\hat{F}_{i}(\omega_{i},d)$ for some $\hat{F}_{i}$, and the guarantee of privacy is presented in Theorem 6.2 (output perturbation) and 6.3 (objective perturbation) below. The total privacy loss in the following theorem is composed using moments accountant method (Abadi et al., 2016).
Theorem 6.2.
Theorem 6.3.
Consider the private Upcycled-FL over $2M$ iterations under objective perturbation with noise $n^{m}_{i}\sim\exp{(-\alpha_{i}^{m}||n_{i}^{m}||_{2})}$. Suppose $\hat{F}_{i}$ is generalized linear model (Iyengar et al., 2019; Bassily et al., 2014)In supervised learning, the sample $d=(x,y)$ corresponds to the feature and label pair. Function $\hat{F}_{i}(\omega,d)$ is generalized linear model if it can be written as a function of $\omega^{T}x$ and $y$. that satisfies $||\nabla\hat{F}_{i}(\omega;d)||<u_{1},~{}|\hat{F}^{\prime\prime}_{i}|\leq u_{2}$. Let feature vectors be normalized such that its norm is no greater than 1, and suppose $u_{2}\leq 0.5|\mathcal{D}_{i}|\mu$ holds. Then the algorithm satisfies $(\varepsilon,0)$-DP for agent $i$ where $\varepsilon=\sum_{m=0}^{M}\frac{2\alpha^{m}_{i}u_{1}\mu+2.8u_{2}}{|\mathcal{D}_{i}|\mu}.$
The assumptions on $\hat{F}_{i}$ are again fairly standard in the literature, see e.g., (Chaudhuri et al., 2011; Zhang and Zhu, 2016; Zhang et al., 2018b). Theorem 6.2 and 6.3 show that the total privacy loss experienced by each agent accumulates over iterations and privacy loss only comes from odd iterations. In contrast, if consider differentially private FedProx, accumulated privacy loss would come from all iterations. Therefore, to achieve the same privacy guarantee, private Upcycled-FL requires much less perturbation per iteration than private FedProx. As a result, accuracy can be improved significantly. Experiments in Section 7 show that Upcycled-FL significantly improves privacy-accuracy trade-off compared to other methods.
@@ 7 Experiments
In this section, we empirically evaluate the performance of Upcycled-FL by combining it with several popular FL methods. We first consider non-private algorithms to examine the convergence (rate) and robustness of Upcycled-FL against statistical/device heterogeneity. Then, we adopt both output and objective perturbation to evaluate the private Upcycled-FL.
7.1 Datasets and Networks
We conduct experiments on both synthetic and real data, as detailed below. More details of each dataset are given in Appendix F.1.
@@ 7.1 Datasets and Networks
Synthetic data. Using the method in Li et al. (2020), we generate Syn(iid), Syn(0,0), Syn(0.5,0.5), Syn(1,1), four synthetic datasets with increasing statistical heterogeneity. We use logistic regression for synthetic data.
Real data. We adopt two real datasets: 1) FEMNIST, a federated version of EMNIST (Cohen et al., 2017). Here, A multilayer perceptron (MLP) consisting of two linear layers with a hidden dimension of 14x14, interconnected by ReLU activation functions, is used to learn from FEMNIST; 2) Sentiment140 (Sent140), a text sentiment analysis task on tweets (Go et al., 2009). In this context, a bidirectional LSTM with 256 hidden dimensions and 300 embedding dimensions is used to train on Sent140 dataset.
7.2 Experimental setup
All experiments are conducted on a server equipped with multiple NVIDIA A5000 GPUs, two AMD EPYC 7313 CPUs, and 256GB memory. The code is implemented with Python 3.8 and PyTorch 1.13.0 on Ubuntu 20.04. We employ SGD as the local optimizer, with a momentum of 0.5, and set the number of local update epochs E to 10 at each iteration $m$. Note that without privacy concerns, any classifier and loss function can be plugged into Upcycled-FL. However, if we adopt objective perturbation as privacy protection, the loss function should also satisfy assumptions in Theorem 6.3. We take the cross-entropy loss as our loss function throughout all experiments.
@@ 7.2 Experimental setup
To simulate device heterogeneity, we randomly select a fraction of devices to train at each round, and assume there are stragglers that cannot train for full rounds; both devices and stragglers are selected by random seed to ensure they are the same for all algorithms.
Baselines. To evaluate the effectiveness of our strategy, we apply our Upcycled-FL method to seven representative methods in federated learning. We use the grid search to find the optimal hyperparameters for all methods to make a fair comparison. These methods include:
  • FedAvg (McMahan et al., 2017): FedAvg learns the global model by averaging the client’s local model.
  • FedAvgM (Hsu et al., 2019): FedAvgM introduces momentum while averaging local models to improve convergence rates and model performance, especially in non-i.i.d. settings.
  • FedProx (Li et al., 2020): FedProx adds a proximal term to the local objective function, which enables more robust convergence when data is non-i.i.d. across different clients.
  • Scaffold (Karimireddy et al., 2020): Scaffold uses control variates to correct the local updates, which helps in dealing with non-i.i.d. data and accelerates convergence.
  • FedDyn (Acar et al., 2021): FedDyn considers a dynamic regularization term to align the local model updates more closely with the global model.
  • pFedMe (T Dinh et al., 2020): pFedMe is a personalization method to handle client heterogeneity. We set the hyperparameter $k$ in pFedMe as 5 to accelerate the training.
  • FedYogi (Reddi et al., 2021): FedYogi considers the adaptive optimization for the global model aggregation.
Unless explicitly stated, the results we report are averaged outcomes over all devices. More details of experimental setup are in Appendix F.1.
Table 1: Average accuracy and standard deviation with 90% straggler on the testing dataset in the non-private setting over four runs: models are trained over synthetic data (Syn) for 80 iterations (160 for upcycled version), trained over Femnist for 150 iterations (300 for upcycled version), and trained over Sent140 for 80 iterations (160 for upcycled version). We use the grid search to find the optimal results for all methods.
Method Dataset
Syn(iid) Syn(0,0) Syn(0.5,0.5) Syn(1,1) FEMNIST Sent140
FedAvg 98.06±0.07 79.28±0.61 81.58±0.43 80.40±1.28 81.38±3.54 76.11±0.11
Upcycled-FedAvg 98.83±0.29 81.46±0.48 82.89±0.17 81.49±0.53 82.10±1.11 76.32±0.45
FedAvgM 98.43±0.07 80.29±0.83 82.60±0.37 80.59±1.28 80.15±3.72 75.7±0.85
Upcycled-FedAvgM 98.72±0.47 81.74±0.42 83.13±0.12 81.37±0.82 81.30±5.23 74.88±2.29
FedProx 96.52±0.07 80.72±0.77 81.99±0.55 81.19±0.19 79.35±0.65 73.94±0.13
Upcycled-FedProx 97.62±0.32 80.88±0.97 83.10±0.83 81.94±0.57 80.33±3.43 74.25±0.34
Scaffold 97.51±0.24 80.26±1.54 82.44±1.66 74.91±2.67 76.83±2.97 76.34±0.56
Upcycled-Scaffold 98.68±0.12 81.10±0.57 82.64±1.39 76.14±1.28 77.88±5.36 77.34±0.22
FedDyn 97.00±0.19 81.62±0.97 80.64±0.81 77.27±2.95 81.76±0.98 75.97±0.35
Upcycled-FedDyn 98.32±0.08 82.41±1.04 82.89±0.80 80.03±3.02 83.33±0.71 76.03±0.58
pFedme 96.30±0.14 89.15±0.24 89.43±0.67 93.06±0.27 71.73±4.30 72.81±0.85
Upcycled-pFedme 96.77±0.12 89.08±0.22 89.55±0.62 93.12±0.23 76.66±3.37 74.07±0.74
FedYogi $99.30_{\pm 0.35}$ $81.20_{\pm 2.64}$ $79.49_{\pm 1.30}$ $78.95_{\pm 1.98}$ $73.53_{\pm 7.73}$ $77.02_{\pm 0.07}$
Upcycled-FedYogi $\textbf{99.41}_{\pm 0.32}$ $\textbf{81.65}_{\pm 2.44}$ $\textbf{80.84}_{\pm 1.30}$ $\textbf{80.17}_{\pm 0.84}$ $\textbf{75.64}_{\pm 3.17}$ $\textbf{77.59}_{\pm 0.26}$
7.3 Results
Convergence and Heterogeneity.